# Can Video Large Multimodal Models Think Like Doubters—or Double-Down: A Study on Defeasible Video Entailment

### v1

@@ 5.1 Counterfactual Thought-Induced Rational Generation
To capture the complexity of logical semantic relationships among the update, premise, and hypothesis, we leverage VLMMs to generate two rationales for each sample—one explaining why the update acts as a Strengthener and the other explaining why it serves as a Weakener. This approach enables counterfactual reasoning (He et al., 2022; Chen et al., 2023a; Weinzierl and Harabagiu, 2024) by examining both perspectives for each case as follows,
@@
$P(H,U)$ is the prompt as shown in Appendix B.4. $R_{w}$ and $R_{s}$ represent the rationales generated by the model, where $R_{w}$ explains why the update weakens the hypothesis, and $R_{s}$ explains why the update strengthens it. Our method prompts the model to analyze the update from two opposing perspectives: (1) assuming it strengthens the hypothesis and (2) assuming it weakens the hypothesis.
@@ 5.2 ASR-Enhanced Video Content Integration
After generating the two rationales, we must determine which one is more plausible as the final answer. However, because vision-based fine-tuning can sometimes impair an LLM’s original generative or reasoning capabilities (Yan et al., 2024; Chai et al., 2024), we rely on a separate LLM to evaluate which rationale is stronger. Since this LLM cannot directly process visual information, its selection process may introduce biases or yield suboptimal results. To address this, we convert the video content into a textual description, enabling the LLM to incorporate the visual context when comparing the two rationales. Specifically, we input the video into VLMMs to generate the textual description, formulated as $D^{\prime}=\operatorname{VLMM}(V,P_{d})$, where $D^{\prime}$ is the generated description. $P_{D}$ is the prompt, as shown in Appendix B.4.
However, existing VLMMs often generate incomplete or inaccurate video descriptions, as they primarily rely on visual inputs and may overlook crucial auditory information (Nishimura et al., 2024). This issue is especially problematic in Defeasible Video Entailment task since the spoken dialogue conveys key semantic details. To address this, we refine video descriptions by integrating VLMM-generated description with ASR transcripts, ensuring a more comprehensive and context-aware representation of the video. Specifically, we first extract spoken dialogue from the video via an ASR engine to capture verbal context, and then employ an LLM to integrate the ASR transcript with the VLMM-generated description as follows,
@@ 6 Improve Generation Ability of VLMMs
As we mentioned before, there are two limitations in VLMMs: 1) missing key audio information and 2) cannot fully capture the entailment semantics between update, hypothesis, and premise. To address the limitations, we proposed an LLM-Guided ASR-Integrated Generation Framework using the modules devised in our classification framework, as shown in Section 3. Specifically, we first refine VLMM-generated descriptions by ASR-Enhanced Video Content Integration, ensuring more accurate textual representations of the video content. This step is identical to the refinement process in the classification task.
@@ 7 Experiment
Table 3: Performance comparison of different VLMMs on the classification task. SRatio stands for the ratio of Strengthener. TV stands for Text and Vision modalities, and TVA stands for the Text, Vision, and Audio modalities. $\uparrow$ indicates that a higher value is better, while $\leftrightarrow$ indicates that values closer to 50 are preferable. For each model, its best results in each metric are highlighted in bold.
@@ 7.1 Experimental Setup
We select six widely used Vision-Language Multimodal Models (VLMMs) as baselines, categorized into two groups: Video-Only Models, which can process video but do not support audio comprehension, including Video-ChatGPT (Maaz et al., 2024a), VideoGPT+ (Maaz et al., 2024b), VideoLLaMA_V (Vision-Language Branch only) (Zhang et al., 2023), and VideoLLaMA2_V (Vision-Language Branch only) (Cheng et al., 2024); and Video-Audio Models, which can process both video and audio, including VideoLLaMA_AV (incorporating both Vision-Language and Audio-Language Branches) and VideoLLaMA2_AV (incorporating both Vision-Language and Audio-Language Branches). The prompts remain consistent across all VLMMs, with details provided in Appendix B.1. For the ASR engine, we use whisper (Radford et al., 2023).
Model w/Ours ROUGE-L $\uparrow$ BLEU-4 $\uparrow$ BERTScore $\uparrow$ C1 $\uparrow$ C2 $\downarrow$ C3 $\downarrow$ C4 $\downarrow$ C5 $\downarrow$ C6 $\downarrow$ C7 $\downarrow$ C8 $\downarrow$
VideoChatGPT 14.77 2.31 86.05 39.29 11.12 17.08 18.65 0.00 7.91 0.36 5.60
17.78 3.04 87.37 64.86 0.83 5.89 23.72 0.00 4.16 0.00 1.05
VideoGPT+ 26.57 6.16 89.11 35.74 4.38 5.99 23.68 0.06 28.91 0.06 1.18
27.31 5.80 89.36 52.78 1.46 7.23 29.24 0.00 7.87 0.00 1.42
VideoLLaMA_V 5.95 0.68 82.09 7.50 29.31 7.94 7.23 1.30 12.31 0.00 34.41
13.66 1.71 86.20 60.63 1.03 5.98 27.20 0.00 4.26 0.03 0.86
VideoLLaMA_AV 6.11 0.77 82.29 13.96 23.06 9.58 9.19 0.92 17.36 0.00 25.93
13.66 1.71 86.20 63.20 1.40 6.02 23.79 0.04 4.77 0.04 0.75
VideoLLaMA2_V 22.59 4.61 87.25 48.68 9.88 15.40 14.77 0.00 8.36 0.00 2.91
21.47 4.03 88.13 56.99 1.41 4.68 29.33 0.00 6.67 0.03 0.90
VideoLLaMA2_AV 22.37 4.73 87.14 47.04 11.53 14.67 14.41 0.00 9.39 0.07 2.89
21.35 4.12 88.10 58.76 1.08 4.21 28.71 0.00 6.33 0.07 0.91
Table 4: Evaluation of VLMMs on the generation task. Each number in C1-C8 (%) indicates the proportion of model-generated updates that fall into the respective category.
@@ 7.1.1 Evaluation Metric
To evaluate the performance of the model $\mathcal{M}_{c}$ on the classification task, we use multiple typical classification metrics: F1, accuracy, recall, and precision.This evaluation framework provides a balanced and detailed view of the model’s classification capabilities across multiple dimensions. To comprehensively assess the quality of the generation model $\mathcal{M}_{g}$, we first utilize traditional string-matching-based metrics ROUGE-L (Lin, 2004) and BLEU-4 (Papineni et al., 2002), and deep learning-based metric such as BERTScore (Zhang et al., 2020). However, these metrics assess the lexical or semantic similarity between the generated update and the reference updates. For open-domain tasks, where acceptable answers can vary widely, assembling a complete set of ground-truth references is impractical.
To address this, we designed a set of carefully defined categories to analyze generated updates, focusing on their interactions with the premise and hypothesis. Specifically, we define eight categories as follows: C1 Good: The update strengthens the hypothesis if the target update type is Strengthener, or weakens the hypothesis if the target update type is Weakener. C2 Neutral: The update neither strengthens nor weakens the hypothesis. C3 Weakener instead of Strengthener: The update, generated as a Strengthener, actually weakens the hypothesis. C4 Strengthener instead of Weakener: The update, generated as a Weakener, actually strengthens the hypothesis. C5 Restating the Premise: The update simply describes what is shown in the video, and does not contribute toward evaluating the hypothesis. C6 Restating the Hypothesis: The update roughly repeats the hypothesis. C7 Contradicting the Premise: The update contradicts the premise, either explicitly or implicitly. C8 Nonsensical or Other: The update itself is nonsensical, incoherent, or unrelated.
These categories enable a detailed assessment of the overall update quality, the specific error patterns, and their alignment with the intended update type. Since human evaluation is costly, we leverage GPT-4o to categorize the generated output into one of the eight predefined categories. Specifically, we provide GPT-4o with the video description from VideoLLaMA2, the hypothesis, and the generated update, allowing it to determine the appropriate category for the output. The prompt is shown in Appendix B.3.
@@ 7.2 Main Results
Classification Task. We show the classification performance comparison across baselines in Table 3. We have several observations. 1) Incorporating audio information enhances model performance, as models with audio input consistently achieve higher F1 scores. For example, the F1 of VideoLLaMA improves from 42.47% to 51.69% with the addition of audio, while VideoLLaMA2 sees a smaller yet notable increase from 65.98% to 67.60%. 2) Newer generations of models demonstrate improved performance, with more recent models such as VideoGPT+ and VideoLLaMA2 outperforming their predecessors, VideoGPT and VideoLLaMA_V. Specifically, VideoGPT+ achieves an F1 of 70.02%, significantly surpassing Video-ChatGPT (45.61%), while VideoLLaMA2_V attains 65.98%, outperforming the earlier VideoLLaMA_V model (42.47%). These results indicate that models with stronger foundational capabilities tend to perform better in our classification task. 3) Our framework consistently improves all baselines across accuracy and F1, demonstrating our method’s effectiveness. Notably, adding our method results in a more balanced prediction distribution, with SRatio values shifting closer to 50%, indicating a more equitable and unbiased classification of strengtheners and weakeners.
Generation Task. Table 4 presents the performance of VLMMs on the generation task. For the traditional metrics BLEU-4, ROUGE-L, and BERTScore, similar to the classification task, successor models outperform their predecessors, with VideoGPT+ surpassing VideoGPT and VideoLLaMA2 achieving higher scores than VideoLLaMA across all three metrics. Our generation framework improves VideoChatGPT, VideoGPT+, VideoLLaMA_V, and VideoLLaMA_AV across most generation metrics. However, these metrics primarily measure lexical and semantic similarity between reference texts and generated texts and are not suitable for our open-domain generation task.
Based on our devised evaluation method, we observed that 1) VideoLLaMA2_V and VideoLLaMA2_AV demonstrate the strongest generation ability for our generation task, as reflected in their high C1 and low C3/C4. 2) In contrast, VideoGPT+ and Video-ChatGPT perform poorly in reasoning precision, as evidenced by their higher C2, C3, and C4, suggesting that they frequently generate neutral updates or confuse Strengtheners with Weakeners. The potential reason may be their low instruction-following ability. 3) Additionally, VideoGPT+ shows a higher C5 and C6, revealing that it tends to restate the premise or hypothesis instead of constructing meaningful updates. 4) Our method significantly improves the proportion of correctly generated updates (C1) across all models, indicating the effectiveness of our generation method. 5) Furthermore, the reduction in C3 suggests that our approach helps models better differentiate between strengthening and weakening updates. 6) Additionally, C8 sees a notable reduction across models, demonstrating improved coherence in generated updates. For example, VideoLLaMA_V reduces nonsensical outputs from 34.41% to 0.86%. In addition, we also add human evaluation in the Appendix A.
@@ 7.3 Ablation Study
To better understand the contributions of different components in our classification framework, we conduct an ablation study by removing each module: (1) ASR-enhanced video descriptions, and (2) Counterfactual Thought reasoning. Table 5 presents the performance impact of these ablations. Without ASR-enhanced descriptions, models tend to rely more on textual artifacts rather than comprehensively understanding the video content. For instance, VideoLLaMA2_AV exhibits a recall drop from 72.21% (w/ ours) to 44.83% (w/o refined description), indicating that without enriched descriptions, models struggle to correctly classify entailment cases that require spoken content understanding. Excluding Counterfactual Thought leads to a significant increase in errors where updates are misclassified as Strengtheners instead of Weakeners (or vice versa). This is reflected in the recall drop in VideoLLaMA2_V from 77.89% to 49.70%, showing that the model without counterfactual reasoning guidance performs worse. These results confirm that our proposed enhancements effectively mitigate VLMMs’ limitations, leading to a more accurate and interpretable classification framework.
@@
Table 5: Ablation study on the classification task, evaluating the effects of ASR-enhanced descriptions, counterfactual reasoning, and rationale refinement. For each model, its best results in each metric are highlighted in bold.
@@ 8 Conclusion
In this work, we introduce Defeasible Video Entailment (DVidE), an extension of the video entailment task that requires dynamically revising inferences based on new updates. To enhance classification performance, we propose the Chain of Counterfactual Thought framework, which leverages counterfactual reasoning, ASR-enhanced Video Content Integration, and Rationale Refinement to reduce inference bias. For generation, we develop a framework that integrates an external ASR tool and LLM to produce coherent, contextually relevant updates aligned with the intended strengthener or weakener goal. Our approach significantly improves performance. Additionally, we introduce the Defeasible Video Entailment benchmark, providing a dataset with strengthener/weakener annotations to support future research. To address the limitations of traditional automatic metrics, we propose a novel LLM-based evaluation method for the generation task.
@@ Appendix A Human Evaluation
To further validate the effectiveness of our method, we conducted a human evaluation, as previous research [Jing et al., 2024] has shown that automatic evaluation may introduce unreliable errors in generation tasks. To optimize cost-efficiency, we randomly sampled 100 instances from two representative models, VideoGPT+ and VideoLLaMA2, including both their original and refined versions with our framework, resulting in a total of 400 samples. Each update was annotated by three evaluators, who classified it into one of the eight predefined categories. The final category assignment was determined by a majority vote among the annotators. The results are presented in Table 6. From Table 6, we observe that our refinement method significantly increases the proportion of high-quality updates (C1), particularly for VideoGPT+, where the percentage of Good updates rises from 21% to 88%. Similarly, VideoLLaMA2 sees an improvement from 50% to 84%. The decrease in C6 suggests that refined models generate fewer redundant updates and are better at introducing meaningful modifications. These results confirm that our approach effectively enhances update generation by improving the logical consistency and informativeness of model outputs. Furthermore, we conduct the case study for qualitative analysis in the Appendix D.

### v2

@@ 5.1 Counterfactual Thought-Induced Rational Generation
To capture the complexity of logical semantic relationships among the update, premise, and hypothesis, we leverage VLMMs to generate two rationales—one explaining why the update acts as a Strengthener and the other explaining why it serves as a Weakener. This approach enables counterfactual reasoning He et al. (2022); Chen et al. (2023a); Weinzierl and Harabagiu (2024) by examining both perspectives for each case as follows,
@@
$P(H,U)$ is the prompt (see Appendix B.4 for details) and $R_{w}$ and $R_{s}$ represent the rationales generated by the model, where $R_{w}$ explains why the update weakens the hypothesis, and $R_{s}$ explains why it strengthens it. Our method prompts the model to analyze the update from two opposing perspectives: (1) assuming it strengthens the hypothesis and (2) assuming it weakens the hypothesis.
@@ 5.2 ASR-Enhanced Video Content Integration
After generating the two rationales, we must determine which one is more plausible as the final answer. However, because vision-based fine-tuning can sometimes impair an LLM’s original generative or reasoning capabilities Yan et al. (2024); Chai et al. (2024), we rely on a separate LLM to evaluate which rationale is stronger. Since this LLM cannot directly process visual information, its selection process may introduce biases or yield suboptimal results. To address this, we convert the video content into a textual description, enabling the LLM to incorporate the visual context when comparing the two rationales. Specifically, we input the video into VLMMs to generate the textual description, formulated as $D^{\prime}=\operatorname{VLMM}(V,P_{d})$, where $D^{\prime}$ is the generated description. $P_{D}$ is the prompt, as shown in Appendix B.4.
However, existing VLMMs often generate incomplete or inaccurate video descriptions, as they primarily rely on visual inputs and may overlook crucial auditory information Nishimura et al. (2024). This issue is especially problematic in Defeasible Video Entailment task since the spoken dialogue conveys key semantic details. To address this, we refine video descriptions by integrating VLMM-generated description with ASR transcripts, ensuring a more comprehensive and context-aware representation of the video. Specifically, we first extract spoken dialogue from the video via an ASR engine to capture verbal context, and then employ an LLM to integrate the ASR transcript with the VLMM-generated description as follows,
@@ 6 Improve Generation Ability of VLMMs
As noted earlier, current VLMMs face two key limitations: (1) they often miss crucial audio information, and (2) they fail to fully capture the entailment semantics among the update, hypothesis, and premise. To address these issues, we introduce an LLM-Guided ASR-Integrated Generation Framework, which builds upon the modules developed in our classification framework (Section 3). Specifically, we first refine the VLMM-generated descriptions through ASR-Enhanced Video Content Integration, producing more accurate textual representations of the video content. This step follows the same refinement procedure used in the classification task.
@@ 7 Experiment
Qwen2.5-VL 71.10 73.61 71.10 70.30 36.71
77.18 78.01 77.18 77.01 36.21
InternVL3 74.65 76.74 74.65 74.14 44.52
76.67 77.36 76.67 76.53 33.57
Table 3: Performance comparison of different VLMMs on the classification task. SRatio stands for the ratio of Strengthener. TV stands for Text and Vision modalities, and TVA stands for the Text, Vision, and Audio modalities. $\uparrow$ indicates that a higher value is better, while $\leftrightarrow$ indicates that values closer to 50 are preferable.
@@ 7.1 Experimental Setup
We select eight widely used VLMMs as baselines, categorized into two groups: Video-Only Models, which can process video but do not support audio comprehension, including Video-ChatGPT Maaz et al. (2024a), VideoGPT+ Maaz et al. (2024b), VideoLLaMA_V (Vision-Language Branch only) Zhang et al. (2023), VideoLLaMA2_V (Vision-Language Branch only) Cheng et al. (2024), Qwen2.5-VL Bai et al. (2025) and internVL3 Zhu et al. (2025); and Video-Audio Models, which can process both video and audio, including VideoLLaMA_AV (incorporating both Vision-Language and Audio-Language Branches) and VideoLLaMA2_AV (incorporating both Vision-Language and Audio-Language Branches). All the VLMMs have 7B parameters. The prompts remain consistent across all VLMMs, with details provided in Appendix B.1. For the ASR engine, we use whisper Radford et al. (2023).
@@ 7.1.1 Evaluation Metric
To evaluate the performance of the model $\mathcal{M}_{c}$ on the classification task, we use multiple typical classification metrics: F1, accuracy, recall, and precision. This evaluation framework provides a balanced and detailed view of the model’s classification capabilities across multiple dimensions.
While traditional string-matching metrics such as ROUGE-L Lin (2004) and BLEU-4 Papineni et al. (2002), as well as embedding-based metrics like BERTScore Zhang et al. (2020), are commonly used to evaluate text generation, they primarily capture lexical or shallow semantic similarity. For open-domain tasks, where acceptable outputs can vary widely and it is impractical to construct exhaustive reference sets, these metrics are inadequate. To address this, we designed a set of carefully defined categories to analyze generated updates, focusing on their interactions with the premise and hypothesis. Specifically, we define eight categories as follows: C1 Good: The update strengthens the hypothesis if the target update type is Strengthener, or weakens the hypothesis if the target update type is Weakener. C2 Neutral: The update neither strengthens nor weakens the hypothesis. C3 Weakener instead of Strengthener: The update, generated as a Strengthener, actually weakens the hypothesis. C4 Strengthener instead of Weakener: The update, generated as a Weakener, actually strengthens the hypothesis. C5 Restating the Premise: The update simply describes what is shown in the video, and does not contribute toward evaluating the hypothesis. C6 Restating the Hypothesis: The update roughly repeats the hypothesis. C7 Contradicting the Premise: The update contradicts the premise, either explicitly or implicitly. C8 Nonsensical or Other: The update itself is nonsensical, incoherent, or unrelated.
Model w/Ours C1 $\uparrow$ C2 $\downarrow$ C3 $\downarrow$ C4 $\downarrow$ C5 $\downarrow$ C6 $\downarrow$ C7 $\downarrow$ C8 $\downarrow$
VideoChatGPT 39.29 11.12 17.08 18.65 0.00 7.91 0.36 5.60
64.86 0.83 5.89 23.72 0.00 4.16 0.00 1.05
VideoGPT+ 35.74 4.38 5.99 23.68 0.06 28.91 0.06 1.18
52.78 1.46 7.23 29.24 0.00 7.87 0.00 1.42
VideoLLaMA_V 7.50 29.31 7.94 7.23 1.30 12.31 0.00 34.41
60.63 1.03 5.98 27.20 0.00 4.26 0.03 0.86
VideoLLaMA_AV 13.96 23.06 9.58 9.19 0.92 17.36 0.00 25.93
63.20 1.40 6.02 23.79 0.04 4.77 0.04 0.75
VideoLLaMA2_V 48.68 9.88 15.40 14.77 0.00 8.36 0.00 2.91
56.99 1.41 4.68 29.33 0.00 6.67 0.03 0.90
VideoLLaMA2_AV 47.04 11.53 14.67 14.41 0.00 9.39 0.07 2.89
58.76 1.08 4.21 28.71 0.00 6.33 0.07 0.91
Qwen2.5-VL 62.17 30.02 0.66 0.10 0.00 0.41 0.56 6.08
83.27 8.32 0.15 0.00 0.00 0.51 0.56 7.20
InternVL3 67.65 23.12 0.86 0.05 0.00 0.46 0.86 7.00
80.27 11.51 0.20 0.00 0.00 0.46 0.61 6.95
Table 4: Evaluation of VLMMs on the generation task. Each number in C1–C8 (%) indicates the proportion of model-generated updates that fall into the category.
These categories give a detailed assessment of the overall update quality, the specific error patterns, and their alignment with the intended update type. Since human evaluation is costly, we leverage GPT-4o to categorize the generated output into one of the eight predefined categories. Specifically, we provide GPT-4o with the video description, the hypothesis, and the generated update, allowing it to determine the appropriate category. The prompt is in Appendix B.3.
@@ 7.2 Main Results
Classification Task. We show the classification performance comparison across baselines in Table 3. We have several observations. 1) Incorporating audio information enhances model performance, as models with audio input consistently achieve higher F1 scores. For example, the F1 of VideoLLaMA improves from 42.47% to 51.69% with the addition of audio, while VideoLLaMA2 sees a smaller yet notable increase from 65.98% to 67.60%. 2) Newer generations of models demonstrate improved performance, with more recent models such as VideoGPT+ and VideoLLaMA2 outperforming their predecessors, VideoGPT and VideoLLaMA_V. Specifically, VideoGPT+ achieves an F1 of 70.02%, significantly surpassing Video-ChatGPT (45.61%), while VideoLLaMA2_V attains 65.98%, outperforming the earlier VideoLLaMA_V model (42.47%). The Latest models Qwen2.5-VL and InternVL3 have the best F1. These indicate that models with stronger foundational capabilities tend to perform better in our classification task. 3) Our framework consistently improves all baselines across accuracy and F1, demonstrating our method’s effectiveness. Notably, adding our method results in a more balanced prediction distribution, with SRatio values shifting closer to 50%, indicating a more unbiased classification of strengtheners and weakeners.
Generation Task. Table 4 presents the performance of VLMMs on the generation task. We observed that 1) Qwen2.5-VL and InternVL3 demonstrate the strongest generation ability for our generation task, as reflected in their high C1 and low C3/C4. 2) In contrast, VideoGPT+ and Video-ChatGPT perform poorly in reasoning precision, as evidenced by their higher C2, C3, and C4, suggesting that they frequently generate neutral updates or confuse Strengtheners with Weakeners. The potential reason may be their low instruction-following ability. 3) VideoGPT+ shows a higher C5 and C6, revealing that it tends to restate the premise or hypothesis instead of constructing meaningful updates. 4) Our method significantly improves the proportion of correctly generated updates (C1) across all models, indicating the effectiveness of our generation method. 5) Furthermore, the reduction in C3 suggests that our approach helps models better differentiate between strengthening and weakening updates. 6) Additionally, C8 sees a notable reduction across models, demonstrating improved coherence in generated updates. For example, VideoLLaMA_V reduces nonsensical outputs from 34.41% to 0.86%. In addition, we also add human evaluation in the Appendix A.
@@ 7.3 Ablation Study
To better understand the contributions of different components in our classification framework, we conduct an ablation study by removing each module: (1) ASR-enhanced video descriptions, and (2) Counterfactual Thought reasoning. Table 5 presents the performance impact of these ablations. Without ASR-enhanced descriptions, models tend to rely more on textual artifacts rather than comprehensively understanding the video content. For instance, VideoLLaMA2_AV exhibits a recall drop from 72.21% (w/ ours) to 44.83% (w/o refined description), indicating that without enriched descriptions, models struggle to correctly classify entailment cases that require spoken content understanding. Excluding Counterfactual Thought leads to a significant increase in errors where updates are misclassified as Strengtheners instead of Weakeners (or vice versa). This is reflected in the recall drop in VideoLLaMA2_V from 77.89% to 49.70%, showing that the model without counterfactual reasoning guidance performs worse. These results confirm that our proposed enhancements effectively mitigate VLMMs’ limitations.
@@
Table 5: Ablation study on the classification task, evaluating the effects of ASR-enhanced descriptions, counterfactual reasoning, and rationale refinement.
@@ 8 Conclusion
We introduce Defeasible Video Entailment, an extension of the video entailment task that requires dynamically revising inferences based on new updates. To enhance classification performance, we propose the Chain of Counterfactual Thought framework, which leverages counterfactual reasoning, ASR-enhanced Video Content Integration, and Rationale Refinement to reduce inference bias. For generation task, we have developed a framework that integrates an external ASR tool and LLM to produce coherent, contextually relevant updates aligned with the intended strengthener or weakener goal. Our approach significantly improves performance. Additionally, we introduce the Defeasible Video Entailment benchmark, providing a dataset with strengthener/weakener annotations to support future research. To address the limitations of traditional automatic metrics, we propose a novel LLM-based evaluation method for the generation task. Our work advances dynamic video reasoning and multimodal inference, laying the foundation for improving VLMMs’ video understanding capabilities.
9 Limitations
@@ Appendix A Human Evaluation
To further validate the effectiveness of our method, we conducted a human evaluation, as previous research (Jing et al., 2024) has shown that automatic evaluation may introduce unreliable errors in generation tasks. To optimize cost-efficiency, we randomly sampled 100 instances from two representative models, VideoGPT+ and VideoLLaMA2, including both their original and refined versions with our framework, resulting in a total of 400 samples. Each update was annotated by three evaluators, who classified it into one of the eight predefined categories. The final category assignment was determined by a majority vote among the annotators. The results are presented in Table 6. From Table 6, we observe that our refinement method significantly increases the proportion of high-quality updates (C1), particularly for VideoGPT+, where the percentage of Good updates rises from 21% to 88%. Similarly, VideoLLaMA2 sees an improvement from 50% to 84%. The decrease in C6 suggests that refined models generate fewer redundant updates and are better at introducing meaningful modifications. These results confirm that our approach effectively enhances update generation by improving the logical consistency and informativeness of model outputs. Furthermore, we conduct the case study for qualitative analysis in the Appendix E.